Find t so that (-8,-10) is a point on the line tx−5y=−7
Answer by htmentor(1343)

Find t so that (-8,-10) is a point on the line tx−5y=−7
=========================
Put the equation in standard form, y = mx + b:
y = (t/5)x + 7/5
The slope of the line = m = t/5
One point on the line is (-8,-10) and the other point is (0,b) = (0,7/5)
m = t/5 = rise/run = (7/5 - -10)/(0 - -8)
Solve for t:
t/5 = (57/5)/8 = 57/40 -> t = 57/8